The Miracle Company had 36,000 units in process on December 31, 2020 which was 80% complete as to materials but only 40% complete as to conversion costs. The company's records show 56,000 units were transferred to the Finished Goods Inventory during January 2021. On January 31, 2021, 31,000 units were on hand which were 20% complete as to conversion costs and 70% complete as to materials. What are the equivalent units of production (EUPs) for the conversion costs in January, assuming Miracle uses first-in, first-out (FIFO)?

Multiple Choice

  • 59,600 EUPs

  • 52,800 EUPs

  • 47,800 EUPs

  • 67,000 EUPs
Answer #1

answer:47800EUPs

Explanation:

   Given unit completed and transferred to finished goods inventory=56000

   Ending inventory= 31000

   EQUIVALENT UNIT OF PRODUCTION FOR CONVERSION COST ONLY (FIFO METHOD)

   UNITS completed and transferred to finished goods

   Beginning inventory= 36000×60%completion = 21600units

   Units transferred = 56000-36000×100%completion= 20000units

  ENDING INVENTORY

   31000× 20%= 6200

Therefore total no of equivalent units of production for conversion cost under fifo method

= 21600units +20000units +6200units = 47800EUPs

ALL THE BEST

PLEASE DO SUPPORT US

ANY DOUBT PLEASE COMMENT BELOW

Thank you